Homework Assignment 1 Due Wednesday October 2

William Stein

Math 124 Harvard University Fall 2002

Instructions: Please work with others, and acknowledge who you work with in your write up. If you can do a problem using a computer please do, but describe how you use the computer. For more practice you can do the problems in the book.

There are six problems.

1. (1 point each) Compute the following gcd’s using the Euclidean algorithm (show the steps):

gcd(7, 19), gcd(388, 32), and gcd(510, 900).

2. (2 points) Use the Euclidean algorithm to find integers x, y ∈ ℤ such that

123x + 567y = 6.

3. (2 points each) Let R = ℤ[\sqrt{−5}] be the ring of elements of the form a + b\sqrt{−5} such that a, b ∈ ℤ. An nonzero non-unit x in R is irreducible if the only divisors of x are of the form xu with u a unit. Also, the norm of a + b\sqrt{−5} is

N(a + b\sqrt{−5}) = (a + b\sqrt{−5})(a − b\sqrt{−5}) = a^{2} + 5b^{2}.

(a) Find the units in R.

(b) Prove that if x, y ∈ R then N(xy) = N(x)N(y).

(c) Show that 2 is irreducible in the ring ℤ[\sqrt{−5}]. [Hint: If 2 = (a + b\sqrt{−5})(c + d\sqrt{−5}) and take norms.]

(d) Show that (1 + \sqrt{−5}) is irreducible in ℤ[\sqrt{−5}].

[Hint: If (1 + \sqrt{−5}) = (a + b\sqrt{−5})(c + d\sqrt{−5}) and take norms.]

4. (4 points) Find the second smallest positive integer x such that

x ≡ 1 (mod 2),
x ≡ 2 (mod 3),
x ≡ 3 (mod 5),
x ≡ 5 (mod 7).

5. (5 points) Suppose that a, b ∈ ℤ/n, and let ã, b̃ ∈ ℤ be lifts of a, b, respectively. Prove that gcd(ã, gcd(b̃, n)) doesn’t depend on the choice of ã, b̃.

6. (a) (2 points) Prove that φ(n) is the number of units in ℤ/n.

(b) (4 points) Prove that φ is multiplicative as follows. Show that the natural map ℤ/mn → ℤ/m × ℤ/n is an injective map of rings, hence bijective by counting, then look at unit groups.